# Phys2Real: Fusing VLM Priors with Interactive Online Adaptation for Uncertainty-Aware Sim-to-Real Manipulation

### v1

@@ IV-B2 Impact of VLM-Only or RMA-Only Physical Parameter Estimation on Policy Performance
To answer Q2, we perform an ablation, removing the adaptation model prediction and conditioning the policy only on the VLM estimate. Conversely, to address Q3, we remove the VLM estimate and only condition the policy on the adaptation model estimate.
@@
When the weight is placed on top (Fig. 4a), the differences become more pronounced. The privileged RMA phase 1 policy has the steepest CDF near the low-error region, while Phys2Real tracks closely behind, showing its ability to handle challenging, unstable CoM dynamics without privileged information. RMA-only and VLM-only baselines diverge further to the right, reflecting large errors and high failure rates. Diffusion policy exhibits the longest tails. We hypothesize this is because during data collection, the human demonstrator can observe the weight location, which biases their actions, but this information is absent at test time.
These results show that Phys2Real consistently shifts both average performance and the full error distribution toward lower values, outperforming baselines and approaching the performance of a policy with privileged ground truth CoM information across both easy and challenging configurations. Additional ablations evaluating robustness to different VLM priors are presented in Appendix A.
@@ IV-C Real-world Evaluation: Hammer Pushing
For hammer pushing, both Phys2Real and DR achieve 100% success across 15 trials, as shown in Table III, demonstrating that the task is solvable for both methods with sufficient training. However, Phys2Real achieves a faster task completion, with an average time of 77.79 s compared to 90.65 s for DR, showing a 14.2% improvement in pushing efficiency. This represents an improvement of approximately 12.9 s per trial, indicating that Phys2Real produces more efficient trajectories and converges to the target pose more quickly. Final position and orientation errors are comparable between the two methods.
V Conclusion
@@ Acknowledgments
The authors thank John Tucker and Rohan Thakker for valuable discussions and feedback. M. Wang is supported by the NASA NSTGRO Fellowship. S. Tian was supported by NSF GRFP Grant No. DGE-1656518.
@@
As the CoM estimate from the VLM deviates further from the ground truth (from $6.0\,\mathrm{cm}$ to $2.0\,\mathrm{cm}$), Phys2Real maintains more consistent performance than Phase 1.5. Most notably, when $\mathrm{CoM}=-2.0\,\mathrm{cm}$}, Phys2Real achieves an 83.3% success rate, compared to only 33.3% for Phase 1.5.
@@ Appendix B VLM CoM Estimation Examples
VLM priors alone may not be physically grounded. While they provide reasonable initial estimates, interactive online adaptation can substantially refine these estimates. For instance, for the T-block with a metal weight at the top, the VLM tends to bias its estimate toward the geometric center, leading the estimate of $\mathrm{CoM}=+4.0\,\mathrm{cm}$} to be lower than the ground truth $\mathrm{CoM}=+6.0\,\mathrm{cm}$}. The standard deviation across $N=5$ images and $M=5$ queries is $0.9\,\mathrm{cm}$.

### v2

@@ IV-B2 Impact of VLM-Only or RMA-Only Physical Parameter Estimation on Policy Performance
For Q2, we perform an ablation without the adaptation model, conditioning the policy only on the VLM estimate. Conversely, for Q3, we remove the VLM estimate and only condition the policy on the adaptation model’s estimate.
@@
When the weight is placed on top (Fig. 4a), the differences become more pronounced. The privileged RMA phase 1 policy has the steepest CDF near the low-error region, while Phys2Real tracks closely behind, showing its ability to handle challenging, unstable CoM dynamics without privileged information. RMA-only and VLM-only baselines diverge further to the right, reflecting large errors and high failure rates. DP exhibits the longest tails. We hypothesize this is because during data collection, the human demonstrator can observe the weight location, which biases their actions, but this information is absent at test time.
These results show that Phys2Real consistently shifts both average performance and the full error distribution toward lower values, outperforming baselines and approaching the performance of a policy with privileged ground truth CoM information across both easy and challenging configurations.
We provide additional ablations evaluating robustness to different VLM priors in Appendix A.
@@ IV-C Real-world Evaluation: Hammer Pushing
For hammer pushing, both Phys2Real and DR achieve 100% success across 15 trials, as shown in Table III, demonstrating that the task is solvable for both methods with sufficient training. However, Phys2Real achieves faster task completion, with an average time of 77.79 s vs. 90.65 s for DR, a 14.2% improvement in task completion time, indicating that Phys2Real produces more efficient trajectories. Final position and orientation errors are comparable between the two methods.
V Future Work
Our work presents many areas of future work, including the following:
Uncertainty-aware multi-parameter estimation. Our work focuses on estimating the center of mass along a single axis. Real-world manipulation requires reasoning about multiple interacting physical properties, including friction, mass, and stiffness. While our framework is designed to accommodate additional physical parameters, the practical challenges of scaling to more physical parameters are left to future work. Furthermore, to account for estimation uncertainty through $\sigma_{\mathrm{vlm}}$, we take the mean of the VLM’s self-reported uncertainty estimates. However, several principled approaches could be used instead. For instance, conformal prediction or next-token prediction probabilities from a multiple-choice formulation (e.g., having the VLM select between discretized CoM ranges) could be explored in future work for better uncertainty calibration.
Broader object and task generalization. We evaluate on two objects (T-block and hammer) with planar pushing. Future work could evaluate on a wider range of object geometries and physical properties, including asymmetrical, deformable, and articulated objects. Additionally, our results demonstrate that the VLM prior is particularly valuable when interaction histories are intermittent, such as in planar pushing, a non-prehensile task with intermittent contact. A future direction is to explore tasks across the contact spectrum: in contact-rich tasks such as tool use or in-hand manipulation, where the adaptation module receives dense signals and the VLM prior may be less critical, and in impulsive tasks such as throwing or striking, where the outcome is determined by a single contact, making the VLM prior more critical.
Overall, Phys2Real demonstrates that foundation models such as VLMs contain useful physical priors that can be refined through interaction. Extending this idea beyond physical parameter estimation, such as world models for vision-language-action (VLA) models, is a promising direction for building robotic systems that combine the semantic knowledge of foundation models with the physical grounding from interacting with the real world.
VI Conclusion
@@ Acknowledgments
The authors thank John Tucker and Rohan Thakker for valuable discussions and feedback.
@@
As the CoM estimate from the VLM deviates further from the ground truth (from $6.0\,\mathrm{cm}$ to $2.0\,\mathrm{cm}$), Phys2Real maintains more consistent performance than Phase 1.5. Notably, when $\mathrm{CoM}=-2.0\,\mathrm{cm}$}, Phys2Real achieves an 83.3% success rate, compared to only 33.3% for Phase 1.5.
@@ Appendix B VLM CoM Estimation Examples
VLM priors alone may not be physically grounded. While they provide reasonable initial estimates, interactive online adaptation can substantially refine these estimates. For instance, for the T-block with a metal weight at the top, the VLM tends to bias its estimate toward the geometric center, leading the estimate of $\mathrm{CoM}=+4.0\,\mathrm{cm}$} to be lower than the ground truth $\mathrm{CoM}=+6.0\,\mathrm{cm}$}. The standard deviation across $V=5$ images and $Q=5$ queries is $0.9\,\mathrm{cm}$.